Monarchism rules dude!

Duc Louis-Philippe d’Orleans

Feb 25, 1808

My fellow parisians,

I am writing to tell you of the coolest spluftiest form of government we can have: Monarchism.

Why is this system so cool? Because under a constitutional monarchy, we have a king, who is a well respected dignitary which the rest of the world can relate to, and sign deals with. We have a leader who is well trained in all that is required to run a country. But we also have a leader who works with and for the people.

The people elect a council of advisors for the king. The king works with them to make the laws, and has veto power over any laws the advisors pass (obviously this may be overriden with a large enough majority). In this way, the king and the people work together to achieve what is best for France.

Sure, there are other experimental forms of government being proposed on the streets of Paris. But only one has withstood the test of time. Salads ARE green and Leafy!! Monarchism is the natural philosophy of the universe.

Monarchism is cool. France is cool. Long live France! Long live the people! Long live the King!